OpenStudy (anonymous):

A 10 ft ladder leans against a wall at an angle theta. The top of the ladder is x feet above the ground. If the bottom of the ladder is pushed toward the wall, find the rate at which x changes with respect to theta when theta = 60 degrees. Express the answer in units feet/degree. Answer in book is = 0.087 ft/deg

OpenStudy (mathmale):

Draw a diagram, labeling everything. If the top of the ladder is x feet above the ground, how is x related to the length of the ladder (10 feet) and to the angle, theta?

OpenStudy (mathmale):

Write a static equation for that (nothing mov

OpenStudy (mathmale):

nothing moving).

OpenStudy (mathmale):

Now differentiate your equation. Your goal is to find d(theta) / dt, where t = time.

OpenStudy (mathmale):

What would you do first?

OpenStudy (anonymous):

\[\sin \theta = x / 10\] I have this.

OpenStudy (anonymous):

now do I multiply both sides by ten and derive?

OpenStudy (anonymous):

The answer in the back of the book is 0.087 ft/deg

OpenStudy (mathmale):

What specifically are you looking for? Describe it in words, please.

OpenStudy (anonymous):

The rate at which x changes with respect to theta. We are assuming that the ladder bottom of the ladder is being pushed towards the wall.

OpenStudy (mathmale):

all right. How would you express that quantity in symbols? It's a derivative.

OpenStudy (anonymous):

\[dx / d\]

OpenStudy (anonymous):

wait dx/ d(theta)

OpenStudy (mathmale):

The rate at which x changes with respect to theta. Which variable depends upon which, in this sit'n?

OpenStudy (mathmale):

The dependent variable goes on top and the ind..t variable on the bottom when you write this rate of change.

OpenStudy (anonymous):

So then we write it as d(theta) / d(x) .

OpenStudy (jdoe0001):

hmm I"d think more like \(\large \left|\cfrac{dx}{d\theta}\right|_{\theta=60^o}\)

OpenStudy (mathmale):

That's the reverse: How fast does the angle change with x? We want, how fast does x change with the angle, theta?

OpenStudy (mathmale):

We need to find the derivative, with respect to theta, of what relationship?

OpenStudy (anonymous):

The relationship between the length of the ladder and x

OpenStudy (mathmale):

Yes, and what is that relationship?

OpenStudy (anonymous):

sin x /10 or csc 10 / x

OpenStudy (mathmale):

But where's theta in those?

OpenStudy (mathmale):

Right: think in terms of a trig function.

OpenStudy (anonymous):

whoops \[\sin \theta = x / 10\] or \[\csc \theta = 10 / x\]

OpenStudy (mathmale):

sin theta = x/10 is correct; the hypo is 10 and x is the height of the top of the ladder abo ve the ground.

OpenStudy (mathmale):

You could certainly solve for theta if you wish, altho it's not necessary. Want to work with sin theta or inverse-sine theta? Your choice.

OpenStudy (anonymous):

sin theta

OpenStudy (mathmale):

sin theta = x/ 10. We want to find the rate of change of x with respect to theta when theta = 60 degrees. Find the derivative, with respect to angle theta, of sin theta = x/10.

OpenStudy (anonymous):

10cos(theta) = x

OpenStudy (mathmale):

but what about the Chain Rule? recall that x supposedly depends upon theta.

OpenStudy (anonymous):

This is where Im stuck. The next section is the chain rule. We haven't covered that yet. But I don't think we can solve this without that rule.

OpenStudy (mathmale):

How would you express "the derivative of x with respect to theta?"

OpenStudy (mathmale):

Use symbols. Evidently you've heard of "chain rule," even tho' you've not formally studied it.

OpenStudy (mathmale):

Find the rate at which x changes with respect to theta when theta = 60 degrees.

OpenStudy (mathmale):

sin theta = x / 10. taking the derivative with respect to theta, we get (1 cos theta = (1/10) (dx/d-theta). Right?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

Now let theta = 60 deg. What is cos theta?

OpenStudy (anonymous):

1/2

OpenStudy (mathmale):

Yes, and so \[\frac{ 1 }{ 2 }=\frac{ 1 }{10}\frac{ dx }{ d }\]

OpenStudy (mathmale):

\[\frac{ 1 }{ 2 }=\frac{ 1 }{10}\frac{ dx }{ d \theta }\]

OpenStudy (mathmale):

Now please solve this for dx/dtheta.

OpenStudy (anonymous):

dx /dtheta = 5

OpenStudy (mathmale):

right. that's 5 what? units of measurement? please refer to the original question.

OpenStudy (anonymous):

5 feet per degree

OpenStudy (mathmale):

exactly. Congrats, you're done!

OpenStudy (mathmale):

Be sure to label your answer with dx/dtheta= ...

OpenStudy (anonymous):

thanks for the help!!!

OpenStudy (mathmale):

Anyquestions? You're welcome!

OpenStudy (anonymous):

yea, that's not the answer in the back of the book. 0.087 ft/deg

OpenStudy (mathmale):

That's a huge discrepancy. Let's go thru this again fast. Is sin theta = x / 10 true?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

Take the deriv of both sides: cos theta * dtheta/dtheta = (1/100 dx/dtheta. True or false?

OpenStudy (mathmale):

False, it's 1/10

OpenStudy (mathmale):

False, it's (1/10) * dx / dtheta.

OpenStudy (mathmale):

Mult both sides by 10 to remove the fraction 1/10.

OpenStudy (anonymous):

yup. So 10 * 1/2 on the left. And dx / dtheta on the right

OpenStudy (mathmale):

Yes. Your result includes the data that theta = 60 degrees. thus, 5= dx / dtheta. Darn. if sin theta = x/10 we could solve for theta: \[\theta=\sin ^{-1}\frac{ x }{ 10 }\]

OpenStudy (mathmale):

Takin gthe deriv. with resp. to theta of both sides, we get what?

OpenStudy (mathmale):

dtheta ------ = 1 dtheta

OpenStudy (mathmale):

So:\[1=\frac{ 1 }{ ^{\sqrt{1-(x/10)^2}} }\frac{ dx }{d \theta }\]

OpenStudy (mathmale):

You'll have to find the value of x when the angle theta is 60 degrees. Sorry, we must be overlooking some important fact here, if our answer is that much different from that in the book.

OpenStudy (anonymous):

Its fine

OpenStudy (mathmale):

I have to get off the 'Net now, but would be glad to continue with you later on. Thanks for your participation and fast responses.

OpenStudy (anonymous):

Thanks for the help!!

OpenStudy (mathmale):

My pleasure!

OpenStudy (anonymous):

I figured out how to get the right answer. The 5 that we got is correct. But you still need to convert that into feet per degree. You do this by multiplying 5 by PI and then dividing by 180.